# Topolectrical circuits – recent experimental advances and developments

### v1

@@ IV.4.1 Nodal knots and topological drumhead surface states
While various experiments have simulated knotted band touchings in 3D parameter space, in both Hermitian and non-Hermitian (exceptional knots) contexts [105, 175, 501], realizing these knots in momentum space have the benefit of accessing their topological surface states at real-space terminations. Electrical circuits prove particularly suitable due to the ease in implementing long-range couplings, which are essential for the construction of nodal knots in momentum space [201, 214, 502, 60, 503, 184, 216, 504, 505, 198]. In general, the topological surface states of nodal knots lie within the interior of the 2-dimensional surface Brillouin zone projections of these knots, and are thus known as “drumhead” surface states [60].
Below, following Ref. [500], we describe the construction of nodal knot models that are suitable for platforms with reciprocal elements i.e. passive RLC circuits. In a reciprocal lattice with two nodes per unit cell, a minimal circuit Laplacian ansatz can be written as
where imposes a uniform offset, is a complex function of the lattice momentum, and and are Pauli matrices. The bands cross in the 3D Brillouin zone where , i.e. when the coefficients of both Pauli matrices disappear. To facilitate the knot construction, is first mapped onto two complex numbers :
such that different knots correspond to different bivariate polynomials in : for instance, the -torus knots correspond to [500]. Such polynomials were already devised for various knots [506, 507, 508], but their embedding onto the 3D Brillouin zone, where are both periodic functions of , have only been considered in the topological nodal context.
As simulated and experimentally demonstrated in Ref. [500], nodal structures can be resolved from the admittance eigenvalues [Fig. 10a]. These nodal knots form intricate patterns in the 3D momentum space, while the corresponding drumhead surface states appear as projected flat bands in the surface admittance spectrum. While the edge-terminated direction was implemented in real space, the two other directions in the plane of the drumhead surface states were parametrically tuned by adjusting the inductance of the inductors to correspond to different momentum values [Fig. 10b]. The inductance was varied using ferrite rods or shorted wire loops to increase or decrease inductance, respectively [Fig. 10c], effectively reduces the undesired effects of parasitics and component uncertainties. By utilizing machine learning-powered optimization and microcontrollers to fine-tune the setup, the drumhead regions could be resolved with a minimal number of measurements [Fig. 10d].
@@ V Machine learning approaches to topolectrical circuit construction and measurement
In addition to data analysis, deep learning can aid the design of the mathematical model itself. Chen et al. [509] utilized multilayer perceptrons and convolutional neural networks to predict topological invariants and the circuit response. They simulated the topolectrical Chern circuit [40] [Fig. 11b1] with 1000 different sets of circuit component parameters to train their model. The topolectrical Chern circuit features chiral voltage edge propagation when a Gaussian pulse is injected at an edge node [Fig. 11b2]. Their trained model demonstrated the approximate phase diagram [Fig. 11b3] and capture aspects of the temporal evolution of chiral voltage propagation, even in the presence of defects [Fig. 11b4].
Chen et al. [510] proposed a framework that integrates text, image, and spectral data to assist in the design of two-dimensional SSH circuits. Their bidirectional collaborative design framework enables both forward prediction of measured quantities from the described circuit structure and inverse design of topolectrical circuits based on textual descriptions. For instance, when provided with an image of a fabricated 2D SSH circuit [Fig. 11c1], the trained model inferred circuit properties that aligned with experimental measurements [Fig. 11c2]. Conversely, given a target voltage profile (e.g., edge-localized voltage [left in Fig. 11c3]), the model suggested a corresponding circuit design [right in Fig. 11c3]. The fabricated circuit based on this predicted structure exhibited a voltage profile similar to the intended design.
@@ VI Outlook
The integration of electrical circuits with advanced computational techniques and emerging technologies [527] promises to deepen our understanding of condensed matter phenomena and push the boundaries of experimental physics [145, 146, 285]. Their adaptability [62, 500] to simulate higher-dimensional systems [50], such as those seen in magic-angle Moiré circuits [528], opens new avenues for investigating uncharted areas, including relatively esoteric condensed matter systems [529, 530, 531, 532, 92, 533, 534]. By continuing to refine and expand the scope of topolectrical circuit implementations, we are well-positioned to uncover fundamental insights and advance the development of practical applications in condensed matter and beyond.

### v2

@@ IV.4.1 Nodal knots and topological drumhead surface states
While various experiments have simulated knotted band touchings in 3D parameter space, in both Hermitian and non-Hermitian (exceptional knots) contexts [105, 175, 516], realizing these knots in momentum space has the benefit of accessing their topological surface states at real-space terminations. Electrical circuits prove particularly suitable due to the ease in implementing long-range couplings, which are essential for the construction of nodal knots in momentum space [202, 215, 517, 60, 518, 184, 217, 519, 520, 199]. In general, the topological surface states of nodal knots lie within the interior of the 2-dimensional surface Brillouin zone projections of these knots, and are thus known as “drumhead” surface states [60].
Below, following Ref. [515], we describe the construction of nodal knot models that are suitable for platforms with reciprocal elements, i.e., passive RLC circuits. In a reciprocal lattice with two nodes per unit cell, a minimal circuit Laplacian ansatz can be written as
where imposes a uniform offset, is a complex function of the lattice momentum, and and are Pauli matrices. The bands cross in the 3D Brillouin zone where , i.e., when the coefficients of both Pauli matrices disappear. To facilitate the knot construction, is first mapped onto two complex numbers :
such that different knots correspond to different bivariate polynomials in : for instance, the -torus knots correspond to [515]. Such polynomials were already devised for various knots [521, 522, 523], but their embedding onto the 3D Brillouin zone, where are both periodic functions of , has only been considered in the topological nodal context.
As simulated and experimentally demonstrated in Ref. [515], nodal structures can be resolved from the admittance eigenvalues [Fig. 10a]. These nodal knots form intricate patterns in the 3D momentum space, while the corresponding drumhead surface states appear as projected flat bands in the surface admittance spectrum. While the edge-terminated direction was implemented in real space, the two other directions in the plane of the drumhead surface states were parametrically tuned by adjusting the inductance of the inductors to correspond to different momentum values [Fig. 10b]. The inductance was varied using ferrite rods or shorted wire loops to increase or decrease inductance, respectively [Fig. 10c], effectively reducing the undesired effects of parasitics and component uncertainties. By utilizing machine learning-powered optimization and microcontrollers to fine-tune the setup, the drumhead regions could be resolved with a minimal number of measurements [Fig. 10d].
@@ V Machine learning approaches to topolectrical circuit construction and measurement
In addition to data analysis, deep learning can aid the design of the mathematical model itself. Chen et al. [524] utilized multilayer perceptrons and convolutional neural networks to predict topological invariants and the circuit response. They simulated the topolectrical Chern circuit [40] [Fig. 11b1] with 1000 different sets of circuit component parameters to train their model. The topolectrical Chern circuit features chiral voltage edge propagation when a Gaussian pulse is injected at an edge node [Fig. 11b2]. Their trained model demonstrated the approximate phase diagram [Fig. 11b3] and captured aspects of the temporal evolution of chiral voltage propagation, even in the presence of defects [Fig. 11b4].
Chen et al. [525] proposed a framework that integrates text, image, and spectral data to assist in the design of two-dimensional SSH circuits. Their bidirectional collaborative design framework enables both forward prediction of measured quantities from the described circuit structure and inverse design of topolectrical circuits based on textual descriptions. For instance, when provided with an image of a fabricated 2D SSH circuit [Fig. 11c1], the trained model inferred circuit properties that aligned with experimental measurements [Fig. 11c2]. Conversely, given a target voltage profile (e.g., edge-localized voltage [Fig. 11c3]), the model suggested a corresponding circuit design [Fig. 11c4]. The fabricated circuit based on this predicted structure exhibited a voltage profile similar to the intended design.
@@ VI Outlook
The integration of electrical circuits with advanced computational techniques and emerging technologies [542] promises to deepen our understanding of condensed matter phenomena and push the boundaries of experimental physics [145, 146, 285]. Their adaptability [62, 515] to simulate higher-dimensional systems [50], such as those seen in magic-angle Moiré circuits [543], opens new avenues for investigating uncharted areas, including relatively esoteric condensed matter systems [544, 545, 546, 547, 92, 548, 549].
Potential future directions in topolectrical circuits include the integration of tunable active elements, such as digitally controlled inductors and analog multipliers, enabling real-time parameter adjustments for Floquet circuit implementations. Although not yet widely available commercially, memristors—often considered the fourth passive circuit element [550, 530, 551] alongside capacitors, inductors, and resistors—could facilitate memory-based physical phenomena such as axon-like transmissions [552] or protected topological states even with a broken symmetry [529]. The incorporation of other nonlinear elements, such as transistors and various types of diodes, remains an open avenue that could simplify circuit implementations and broaden potential applications. These advancements may transition topolectrical circuits from proof-of-concept research to real-world applications in signal processing, computation, and sensing. Further refinement and expansion of the scope of topolectrical circuits could lead to uncovering fundamental insights and advance the development of practical applications in condensed matter and beyond.